OpenStudy (anonymous):

A right triangle has one side of length 8 and a hypotenuse of length 10. How long is the third side of the triangle?

OpenStudy (anonymous):

h=sqrt(x^2+y^2) here h=10 and x=8 put the values in eqn here x and y are sides of tringle and h is hypotenuse so, 10=sqrt(8^2+y^2) so squaring on both sides u wil get  now find value of y..by taking sqr root of y..